OpenStudy (anonymous):

solve for y. 5x-8y=1

jimthompson5910 (jim_thompson5910):

Similar Problem: \[\Large 3x-8y=13\] \[\Large -8y=13-3x\] \[\Large -8y=-3x+13\] \[\Large y=\frac{-3x+13}{-8}\] \[\Large y=\frac{-3x}{-8}+\frac{13}{-8}\] \[\Large y=\frac{3}{8}x-\frac{13}{8}\]

jimthompson5910 (jim_thompson5910):

To solve your problem, use the same steps/idea, but just use different numbers

OpenStudy (anonymous):

Great, thanks!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

mathematically this equation having infinite number of solution....